# PEPT: Expert Finding Meets Personalized Pre-training

### v1

@@ 4. PROPOSED METHOD
In this part, we provide an introduction to our method PEPT. Our method including two stages: pre-training expert representations and fine-tuning with the expert finding task. These stages are illustrated in the Fig 1 and Fig 2.
@@
Figure 1. PEPT Personalized Pre-training Framework. We unify heterogeneous information into the pre-training model to provide rich fundamental information for learning expert representations. During pre-training, we additionally design a question-level masked language model and a vote-oriented task to empower the model abilities to capture expert interest and expertise.
@@ 4.1. Pre-training Stage
The main difference between this work and the conference version lies in the design of the pre-training architecture. In this work, we introduce personalized expert information into the model through the soft prompt to achieve personalized pre-training. Via adopting this approach, the model becomes capable of capturing and understanding the distinct attributes of each individual expert. Additionally, by designing specific question-level pre-training tasks, the model gains the ability to capture more fine-grained expert interests and capabilities, when facing with different target questions.
In the following sections, we present the PEPT from three perspectives: (1) Input layer for introducing the input information constructed for the model; (2) Model architecture for introducing the model components; (3) Pre-training task for introducing our specially tailored training tasks for expert finding.
@@ 4.1.1. Input Layer
Different from traditional language knowledge pre-training in Nature Language Progressing, which concentrates on corpus-level information input, we aim to empower the BERT abilities to personalize capture expert representations. We introduce the input layer from the three separate aspects (i.e., interest, expertise, and personalized).
@@
In CQA websites, the user’s interests can be reflected from the questions he/she has answered historically. And hence, as shown in Fig 1, we construct a unified expert-level sequence by directly concatenating the words from the expert’s previously answered questions, allowing the model to capture the expert’s interest information. Besides, experts’ representations should be different when faced with different target questions. Therefore, we further incorporate the target question information with the expert-level interest information.
@@
Subsequently, taking into account the extensive linguistic knowledge already captured by the original pre-trained BERT weight (Devlin et al., 2018), we employ that to initialize the model parameters. Noted that the segment ID is always to 00, since we don’t employ the Next Sentence Prediction (NSP) task. And the input representation matrix $\mathbf{R}_{t}\in\mathcal{R}^{t_{num}\times d}$ ($t_{num}$ is the number of tokens) is generated by summing its corresponding embeddings:
@@
Specifically, we pre-process vote scores (introduced in Implementation Details) of questions and employ an embedding matrix to map each vote score as an embedding vector (with $d$ as dimension). To accurately capture the expert’s expertise in a specific question, we ensure that the dimension of each vote score embedding matches the dimensionality of the corresponding historical question. In other words, the length of one question token is $x$, and the corresponding vote score embedding matrix is $x\times d$. Then we stack these vote score embeddings as a matrix $\mathbf{R}_{v}\in\mathcal{R}^{t_{num}\times d}$. Afterwards, we integrate $\mathbf{R}_{v}$ with the input representation $\mathbf{R}_{t}$ as follows:
@@
Different from NLP, expert finding further emphasize personalization. User ID embeddings are commonly used in recommendation models and could be considered as the identifying information of individual users (Rendle et al., 2009). Hence, we introduce the expert ID into the pre-training, which could reflect unique expert characteristics and help personalized pre-training. A simple and intuitive approach is to consider expert IDs as distinct word tokens and incorporate them into the pre-trained model’s vocabulary. However, considering potentially millions of experts on the CQA websites, this would lead to an excessively large word vocabulary, making pre-training time-consuming. It’s worth noting that the expert ID doesn’t need to be in word form and can be represented by a vector embedding created through random initialization or other techniques. This type of processing method can be formally termed soft prompt. Therefore, we treat expert IDs as additional prompting tokens rather than adding them to the vocabulary.
@@
where $\mathbf{c}^{u}_{i}\in\mathcal{R}^{1\times d}$, and $f(i)\in{\{1,0\}}^{M}$ represents a one-hot vector, where the $1$ indicates the position of the expert’s vector in $\mathbf{C}^{u}$. Then, we concatenate the ID embedding of the expert at the begin of the corresponding $\mathbf{R}_{in}$, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$ with $\mathcal{R}^{(t_{num}+1)\times d}$ as dimension. This approach provides the foundation for the model to personalized pre-train expert interest and expertise across various questions.
@@ 4.1.2. Model Architecture
where $\mathbf{W}^{q}_{i},\mathbf{W}^{k}_{i},\mathbf{W}^{v}_{i}\in\mathcal{R}^{d\times\frac{d}{h}}$ and $\mathbf{W}\in\mathcal{R}^{d\times d}$ are learnable parameters. Through this function, the $\mathbf{R}_{in}$ is transformed to $\mathbf{H}\in\mathcal{R}^{t_{num}\times d}$ ($t_{num}$ is the token number), which capture the contextual relationships.
@@
Figure 2. PEPT Personalized Fine-tuning. It is noted that [+39] and [-4] represent different question vote scores. The pre-trained encoder is the pre-trained expert weight from PEPT Personalized Pre-training.
@@ 4.1.3. Pre-training Task
We introduce our designed pre-training tasks in this part, which are the core mechanisms that encourage PLMs to model experts’ interest and expertise across various questions.
@@
(9) $\mathcal{L}_{ql}=-logp(\texttt{Jews}|x_{11},x_{16},p_{14}),\texttt{Jews}\in[1,2,\cdots,\lvert\mathcal{V}\rvert]\ ,$
@@
As denoted above, the vote score the expert received could reflect the expertise in answering one question. Generally, higher vote scores indicate higher expertise, and the answer would be more satisfied the CQA users. Therefore, we design a vote-oriented task to pre-train the model and enhance its capacity to discern the expert’s expertise in addressing the target question. Specifically, given the instance shown in Fig 1, the model output embedding from the expert ID can be viewed as an expert-level representation $\mathbf{e}\in\mathcal{R}^{1\times d}$. Hence, given a target question, we directly employ $\mathbf{e}$ to predict the vote score that the expert would receive, which is defined as:
(10) $S_{v}=\mathbf{W}_{v}(\mathbf{e^{T}})+\mu_{v}\ ,$
where $\mathbf{W}_{v}\in\mathcal{R}^{1\times d}$ is the model parameter and $\mu_{v}$ denotes the bias. As the vote prediction is a regression problem, we employ the square loss function to predict the received vote score for pre-training our model, which is formulated as follows:
(11) $\mathcal{L}_{vs}=\sum_{\mathbf{\Omega}}(\hat{S}_{v}-S_{v})^{2}\ ,$
where $\mathbf{\Omega}$ denotes the all instance in the pre-training stage, and $\hat{S}_{v}$ is the ground truth vote score.
Additionally, acknowledging the potent capacity of the original MLM task (Devlin et al., 2018) to train PLMs, we incorporate this task to acquire linguistic knowledge in CQA websites. The MLM task randomly masks certain words and subsequently leverages bidirectional contextual information to reconstruct the input sequence. The formal definition is:
(12) $\mathcal{L}_{wm}=-\sum_{i=1}^{\mathcal{N}}logp(n=n_{i}|\theta),n_{i}\in[1,2,\cdots,\lvert\mathcal{V}\rvert],$
where $\mathcal{N}$ is the masked token number, and $\mathcal{V}$ is the vocabulary.
@@ 4.1.4. Model Learning
During pre-training, the PEPT model is collectively trained in three tasks, and the loss function used is:
@@ 4.2. Fine-tuning Stage
In this section, we perform the expert finding task using a fine-tuning approach. As depicted in Fig 2, we combine the previously answered questions by the expert with the target question and input the resulting sequence directly into the pre-trained model, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$. By adopting this approach, we ensure that the inputs during the fine-tuning stage remain consistent with those used during the pre-training phase. This consistency helps minimize the disparity between the pre-training and fine-tuning stages.
@@
During fine-tuning, we follow previous works (Li et al., 2019; Peng et al., 2022a) and utilize the negative sampling technology (Huang et al., 2013) to generate $k$ negative samples. Then, we employ the cross-entropy loss for optimizing, which can be expressed as follows:
@@ 5. EXPERIMENTAL SETTINGS
To evaluate the performance of PEPT, we conduct a series of comprehensive experiments on six CQA datasets. To begin, we provide an introduction to the datasets employed in our paper. Afterwards, we give the introduction to the implementation details of our experiment. Then, we introduce the baseline methods and evaluation metrics used in this paper.
@@ 5.1. Datasets
For our experiments, we use a dataset comprising 103,005 experts extracted from StackExchange [^2] https://archive.org/details/stackexchange. Each expert corresponds to his/her historical answered questions. This dataset is used for pre-training expert representations in PEPT. We evaluate the effectiveness of the pre-trained model in specialized domains by selecting six distinct domains: English, Biology, Es, Electronics, Gis, and CodeReview. Within each specific domain, we organize a collection of questions. Each question is accompanied by a sequence of title words that provide question semantic information. Additionally, each question has an “accepted answer” chosen from multiple answers provided by different experts. In the evaluation, we consider the expert who has provided the “accepted answer” as the ground truth. To construct candidate experts for questions, we follow the pre-processing method employed in previous research (Peng et al., 2022a). This set includes the original experts who previously answered the question, and randomly selected additional experts from the pool of answerers (one set contains 20 experts in total). Table 3 provides an overview of the statistical properties of the datasets, offering a summary of their key characteristics. To ensure a comprehensive evaluation, we split each dataset into three distinct subsets: training/validation/testing sets. The proportions of these subsets, in chronological order, are set to 80%, 10%, and 10% of the dataset, respectively. Noted that the validation set and the testing set do not appear in the pre-training data.
@@ 5.2. Implementation Details
Within this section, we present a comprehensive overview of the experimental settings and hyperparameters employed in our model. Next, we provide an introduction of the vote score pre-processing technique. It should be noted that all hyperparameters in our model were fine-tuned using the validation dataset.
@@
Our personalized pre-training (PEPT) model, with approximately 110 million parameters, is based on the bert-base-uncased (Devlin et al., 2018). The batch size for both pre-training and fine-tuning is $16$. For alleviating over-fitting, we use dropout technology (Srivastava et al., 2014). Specifically, the dropout ratio is $0.2$ in pre-training and $0.1$ in fine-tuning. We optimize our model using the Adam optimization strategy (Kingma and Ba, 2015). We set the learning rate to $55e-$ 5 during pre-training and $71e-$ 7 during fine-tuning. The token embedding and expert representation dimensions are both $768$. The maximum training step during pre-training is 50,000. We conduct each experiment independently and repeated it three times, and report the average results. All experiments were carried out using the PyTorch framework, which served as the foundation for our implementation. We utilize a single GPU server with an Intel(R) Xeon(R) CPU@2.20GHz and two RTX-3090 GPUs, each with 24GB of memory.
@@
In this part, we introduce how to process the vote scores in our dataset. One of the key challenges we encountered is a large variance in the distribution of vote scores. For example, in the Biology dataset, we observe the maximum vote score is $+287$, while the minimum is $-8$, and the average vote score is $3.49$.
Therefore, to alleviate this situation, we employ a normalization approach. First, we translate the vote scores overall to eliminate negative values (i.e., all vote scores are added to the minimum value in the vote score set). Subsequently, we apply a logarithmic transformation (i.e., $ln(\cdot)$) to the vote scores to reduce the impact of excessive variance. This helps ensure that the scores are more evenly distributed across the dataset. To simplify data processing and ensure roughly equivalent numbers of scores within each score segment, we normalized the vote scores to integers ranging from $1$ to $5$. The normalization process is conducted as follows:
(16) $v_{min}=min(V^{u}),v_{max}=max(V^{u})\ ,$
where $v_{min}$ and $v_{max}$ represent the minimum score and the maximum score of the $V^{u}$ respectively.
(17) $v^{*}_{i}=round(\frac{4}{v_{max}-v_{min}}*(v_{i}-v_{min})+1)\ ,$
here, $v^{*}_{i}$ represents the normalized score of the $v_{i}$ in $V^{u}$, and $round$ denotes the rounding operator.
@@ 5.3. Comparison Methods
To ensure a comprehensive evaluation of the effectiveness of PEPT, we compare its performance against two different types of competitive baselines:
@@
  • Doc2Vec utilizes the Doc2Vec technology to learn expert/question representation and identifies suitable experts based on their past experience in answering similar questions to the target question.
@@
  • EFHM (Peng et al., 2022b) is a multi-grained hierarchical matching framework for expert finding in CQA websites, which could capture more fine-grained matching relationships and achieve superior performance.
@@
Second, the latest PLM-based methods model expert or question features using pre-trained language models, which outperforms above baselines.
@@
Some papers have already published their source code on Github, such as NeRank (Li et al., 2019), PMEF (Peng et al., 2022a), etc. For these models, we download the code and check the model details and parameters based on their original paper to ensure that the model is consistent with the original paper. For models without publicly available source code, such as RMRN (Fu et al., 2020), TCQR (Zhang et al., 2020), etc. We strictly follow technical details reported by the paper to reproduce the model and set model hyperparameters based on the hyperparameter settings reported in the paper.

### v2

@@ 4. PROPOSED METHOD
In this part, we introduce our method PEPT. Our method including two stages: pre-training expert representations and fine-tuning with the expert finding task. These stages are illustrated in the Fig 1 and Fig 2.
@@
Figure 1. PEPT Personalized Pre-training Framework. We unify heterogeneous information into the pre-training model to provide rich information for learning expert representations. During pre-training, we additionally design question-level masked language model and vote-oriented tasks to empower the model abilities to capture expert interest and expertise.
@@ 4.1. Pre-training Stage
The main difference between this work and the conference version is the design of the pre-training architecture. In this work, we introduce personalized expert information into the model through the soft prompt to achieve personalized pre-training. Via adopting this approach, the model can be capable of capturing and the distinct characteristics of each expert. Additionally, by designing specific question-level pre-training tasks, the model gains the ability to capture more fine-grained expert interests and capabilities.
In the following sections, we introduce the PEPT from three perspectives: (1) Input layer, for introducing the input information constructed for the model; (2) Model architecture, for introducing the model components; (3) Pre-training task, for introducing our specially tailored training tasks for expert finding.
@@ 4.1.1. Input Layer
Different from traditional language knowledge pre-training in Nature Language Processing, which concentrates on corpus-level information input, we aim to empower the BERT abilities to personalized capture expert representations. We introduce the input layer from the three separate aspects (i.e., interest, expertise, and personalized).
@@
In CQA websites, the user’s interests can be reflected from the questions he/she has answered historically. And hence, as shown in Fig 1, we construct a unified expert-level sequence by directly concatenating the words from the expert’s previously answered questions, allowing the model to capture the expert’s interest information. Besides, experts’ representations should be different when faced with different target questions. Therefore, we further incorporate the target question with the expert-level interest information.
@@
Subsequently, taking into account the extensive linguistic knowledge already captured by the original pre-trained BERT weight (Devlin et al., 2018), we employ the bert-base-uncased to initialize the model parameters. And the input representation matrix $\mathbf{R}_{t}\in\mathcal{R}^{t_{num}\times d}$ ($t_{num}$ is the number of tokens) is generated by summing its corresponding embeddings:
@@
Specifically, we pre-process vote scores (introduced in Implementation Details) of questions and employ an embedding matrix to map each vote score as an embedding vector (with $d$ as dimension). To accurately capture the expert’s expertise in a specific question, we ensure that the dimension of each vote score embedding matches the dimensionality of the corresponding historical question. In other words, the length of one question token is $x$, and the dimension of corresponding vote score embedding matrix is $x\times d$. Then we stack these vote score embeddings as a matrix $\mathbf{R}_{v}\in\mathcal{R}^{t_{num}\times d}$. Afterwards, we integrate $\mathbf{R}_{v}$ with the input representation $\mathbf{R}_{t}$ as follows:
@@
Different from NLP, expert finding further emphasize personalization. User ID embeddings are commonly used in recommendation models and could be considered as the identifying information of individual users (Rendle et al., 2009). Hence, we introduce the expert ID into the pre-training, which could reflect unique expert characteristics and help personalized pre-training. A simple and intuitive approach is to consider expert IDs as distinct word tokens and incorporate them into the pre-trained model’s vocabulary. However, considering potentially millions of experts on the CQA websites, this would lead to an excessively large word vocabulary, making pre-training time-consuming. It’s worth noting that the expert ID doesn’t need to be in word form and can be represented by a vector embedding created through random initialization or other techniques. Therefore, we treat expert IDs as additional prompting tokens rather than adding them to the vocabulary.
@@
where $\mathbf{c}^{u}_{i}\in\mathcal{R}^{1\times d}$, and $f(i)\in{\{1,0\}}^{M}$ represents a one-hot vector, where the $1$ indicates the position of the expert’s vector in $\mathbf{C}^{u}$. Then, we concatenate the ID embedding of the expert at the begin of the corresponding $\mathbf{R}_{in}$, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$ with $\mathcal{R}^{(t_{num}+1)\times d}$ as dimension. This could prompt the model to personalized pre-train expert interest and expertise across various questions.
@@ 4.1.2. Model Architecture
where $\mathbf{W}^{q}_{i},\mathbf{W}^{k}_{i},\mathbf{W}^{v}_{i}\in\mathcal{R}^{d\times\frac{d}{h}}$ and $\mathbf{W}\in\mathcal{R}^{d\times d}$ are learnable parameters. Through this function, the output is $\mathbf{H}\in\mathcal{R}^{(t_{num}+1)\times d}$ ($t_{num}$ is the token number), which capture the contextual relationships.
@@
Figure 2. PEPT Personalized Fine-tuning. The pre-trained encoder is the pre-trained expert weight from PEPT Personalized Pre-training.
@@ 4.1.3. Pre-training Task
We introduce our designed pre-training tasks in this part, which are the core mechanisms that encourage PLMs to model experts’ personalized interest and expertise.
@@
(9) $\mathcal{L}_{ql}=-logp(\texttt{Jews}|x_{11},x_{16},p_{14}),\texttt{Jews}\in[1,2,\cdots,\lvert\mathcal{V}_{w}\rvert]\ ,$
@@
As denoted above, the vote score the expert received could reflect the expertise in answering one question. Generally, higher vote scores indicate higher expertise, and the answer would be more satisfied the CQA users. Therefore, we design a vote-oriented task to pre-train the model and enhance its capacity to discern the expert’s expertise in addressing the target question. Specifically, given the instance shown in Fig 1, the model output embedding from the expert ID can be viewed as an expert-level representation $\mathbf{e}\in\mathcal{R}^{1\times d}$. Hence, given a target question, we directly employ $\mathbf{e}$ to predict the vote score that the expert would receive, and the loss function is:
(10) $\mathcal{L}_{vs}=-\sum_{i=1}^{\mathcal{S}}logp(v=v_{i}|\theta),v_{i}\in[1,2,\cdots,\lvert\mathcal{V}_{s}\rvert],$
where $\mathcal{S}$ is the masked target question score, and $\mathcal{V}_{s}$ is the vote score vocabulary.
Additionally, since the potent capacity of the original MLM task (Devlin et al., 2018) to train PLMs, we incorporate this task to acquire linguistic knowledge in CQA websites. The MLM task randomly masks certain words and subsequently leverages bidirectional contextual information to reconstruct the input sequence. The formal definition is:
(11) $\mathcal{L}_{wm}=-\sum_{i=1}^{\mathcal{N}}logp(n=n_{i}|\theta),n_{i}\in[1,2,\cdots,\lvert\mathcal{V}_{w}\rvert],$
where $\mathcal{N}$ is the masked token number, and $\mathcal{V}_{w}$ is the word vocabulary.
@@ 4.1.4. Model Learning
During pre-training, the PEPT model is trained with three tasks, and the loss function used is:
@@ 4.2. Fine-tuning Stage
In this section, we perform the expert finding task using a fine-tuning approach. As depicted in Fig 2, we combine the previously answered questions by the expert with the target question and input the sequence directly into the pre-trained model, which is denoted as $[\mathbf{c}^{u}_{i};\mathbf{R}_{in}]$. By adopting this approach, we ensure that the inputs during the fine-tuning stage remain consistent with those used during the pre-training phase. This consistency helps minimize the disparity between the pre-training and fine-tuning stages.
@@
During fine-tuning, we follow previous works (Li et al., 2019; Peng et al., 2022a) and utilize the negative sampling technology (Huang et al., 2013) to generate $k$ negative samples. Then, we employ the cross-entropy loss for optimizing, which can be denoted as:
@@ 5. EXPERIMENTAL SETTINGS
To evaluate the performance of PEPT, we conduct experiments on six CQA datasets. First, we introduce the datasets employed in our paper. Afterwards, we introduce the implementation details of our experiment. Then, we introduce the baseline methods and evaluation metrics used in this paper.
@@ 5.1. Datasets
In experiments, we use a dataset comprising 326781 target-aware expert input information extracted from StackExchange [^5] https://archive.org/details/stackexchange. Each expert corresponds to his/her historical answered questions. This is used for pre-training expert representations in PEPT. We evaluate the effectiveness of the pre-trained model in six domains: English, Biology, Es, Electronics, Gis, and CodeReview. Each question is accompanied by a sequence of title words that provide question semantic information. Additionally, each question has an “accepted answer” chosen from multiple answers provided by different experts. The negative sample in training set is set to 9. In evaluation, the expert who has provided the “accepted answer” is the ground truth. We follow the pre-processing method employed in previous research (Peng et al., 2022a). The candidate expert set for question in evaluation includes the original answerers who previously answered the question, and select additional answerers from the pool of answerers (one set contains 20 experts in total). It is noted that the number of negative samples are same in different methods for fair comparison. Table 3 provides an overview of the statistical properties of the datasets, offering a summary of their key characteristics. We split each dataset into three distinct subsets in chronological order: training, validation and testing sets with 80%, 10% and 10% of the dataset, respectively. Noted that the validation set and the testing set do not appear during pre-training.
@@ 5.2. Implementation Details
In this section, we introduce the experimental settings and hyperparameters employed in our model. Next, we introduce the vote score pre-processing. It is noted that all hyperparameters in our model were fine-tuned using the validation dataset.
@@
Our personalized pre-training (PEPT) model is based on the bert-base-uncased (Devlin et al., 2018). The batch size for both pre-training and fine-tuning is $16$. For alleviating over-fitting, we use dropout technology (Srivastava et al., 2014). Specifically, the dropout ratio is $0.2$ in pre-training and $0.3$ in fine-tuning. We optimize our model using the Adam optimization strategy (Kingma and Ba, 2015). We set the learning rate to $55e-$ 5 during pre-training and $65e-$ 6 during fine-tuning. The token embedding and expert representation dimensions are both $768$. The maximum training step during pre-training is 50,000. We conduct each experiment independently and repeat it three times, and report the average results. All experiments are carried out using the PyTorch framework. We utilize a single GPU server with an Intel(R) Xeon(R) CPU@2.20GHz and two RTX-3090 GPUs, each with 24GB of memory.
@@
In this part, we introduce how to process the vote scores of the questions. One of the key challenges we encountered is a large variance in the distribution of vote scores. For example, in the Biology dataset, we observe the maximum obtained vote score of one question is $+287$, while the minimum is $-8$, and the average vote score is $3.49$.
Therefore, to alleviate this, we employ a normalization approach on a per-question basis. First, we translate the all vote scores to eliminate negative values (i.e., all vote scores are added to the minimum value). Subsequently, we apply a logarithmic transformation (i.e., $ln(\cdot)$) to the vote scores of questions to reduce the impact of excessive variance. This helps ensure that the obtained scores of questions are more evenly distributed across the dataset. To simplify data processing and ensure roughly equivalent numbers of scores within each score segment, we normalized the vote scores to integers ranging from 00 to $9$. The normalization process is conducted as follows:
(15) $v_{min}=min(V),v_{max}=max(V)\ ,$
where $v_{min}$ and $v_{max}$ represent the minimum score and the maximum score of the $V$ (all question vote scores) respectively.
(16) $v^{*}_{i}=round(\frac{9}{v_{max}-v_{min}}*(v_{i}-v_{min})+1)\ ,$
here, $v^{*}_{i}$ represents the normalized score of the $v_{i}$ in $V$, and $round$ denotes the rounding operator.
@@ 5.3. Comparison Methods
We compare the PEPT with two different types of baselines:
@@
  • EFHM (Peng et al., 2022b) is a multi-grained (word-level, question-level and expert-level) hierarchical matching framework for expert finding in CQA websites, which could capture more fine-grained matching relationships and achieve superior performance.
@@
Second, the latest PLM-based methods model expert or question features using pre-trained language models.
@@
Some papers have already published their source code on Github, such as NeRank (Li et al., 2019), PMEF (Peng et al., 2022a), etc. For these models, we download the code and check the model details and parameters based on their original paper to ensure that the model is consistent with the original paper. For models without publicly available source code, such as RMRN (Fu et al., 2020), TCQR (Zhang et al., 2020), etc. We follow technical details reported by the paper to reproduce the model and set model hyperparameters based on the hyperparameter settings reported in the paper.